An aqueous solution of 0.23 M ammonia (NH3) has a pOH of 2.70. Determine the base-dissociation constant (Kb) of ammonia.
Answer #1

Consider reaction, NH3 (aq) + H2O (l) 1585624519074_blob.png NH4 + (aq) + OH - (aq)

Equilibrium constant for above reaction is K b = [ NH4 + ] [ OH - ] / [ NH3 ]

According to above reaction,[ NH4 + ] = [ OH - ]

We have relation, pOH = - log [ OH - ]

[ OH - ] = 10 -pOH  = 10 - 2.70 = 1.995 1585624467753_image.png 10 -03 M

Let's use ICE table.

Concentration ( M) NH3 NH4 + OH -
I 0.23
C - 1.995 1585624467753_image.png 10 -03 + 1.995 1585624467753_image.png 10 -03 + 1.995 1585624467753_image.png 10 -03
E 0.23 - 1.995 1585624467753_image.png 10 -03 1.995 1585624467753_image.png 10 -03 1.995 1585624467753_image.png 10 -03

Kb = ( 1.995 1585624467753_image.png 10 -03 ) ( 1.995 1585624467753_image.png 10 -03 ) / 0.23 - 1.995 1585624467753_image.png 10 -03

Kb = 1.7455 1585624467753_image.png 10 -05

Kb = 1.75 1585624467753_image.png 10 -05